# Improving Socratic Question Generation using Data Augmentation and Preference Optimization

### v1

@@ Abstract
The Socratic method is a way of guiding students toward solving a problem independently without directly revealing the solution to the problem. Although this method has been shown to significantly improve student learning outcomes, it remains a complex labor-intensive task for instructors. Large language models (LLMs) can be used to augment human effort by automatically generating Socratic questions for students. However, existing methods that involve prompting these LLMs sometimes produce invalid outputs, e.g., those that directly reveal the solution to the problem or provide irrelevant or premature questions. To alleviate this problem, inspired by reinforcement learning with AI feedback (RLAIF), we first propose a data augmentation method to enrich existing Socratic questioning datasets with questions that are invalid in specific ways. Next, we propose a method to optimize open-source LLMs such as LLama 2 to prefer ground-truth questions over generated invalid ones, using direct preference optimization (DPO). Our experiments on a Socratic questions dataset for student code debugging show that a DPO-optimized 7B LLama 2 model can effectively avoid generating invalid questions, and as a result, outperforms existing state-of-the-art prompting methods.
Keywords:
Large Language Models Programming Education Socratic Questioning
@@ 1 Introduction
Learning based on a conversation that consists of questions and answers, where the student responds to questions posed by a more knowledgeable instructor, has been proven to be effective in teaching students about a particular concept [35]. In particular, Socratic questioning, which refers to a way for the instructor to guide a student to solve a problem (possibly beyond their zone of proximal development) by asking them questions that promote thinking while not directly revealing the solution [21], is a very relevant pedagogical method in conversation-based learning and tutoring.
@@
Recent advances in large language models (LLMs) [4] have led to the rapid development of chatbots that promote student learning by automatically generating the instructor’s utterances [6, 13, 29]. One key area of interest in the development of such chatbots is question generation, which can help students solve logical problems in the mathematics and programming domains [2, 27]. Typically, question generation in educational applications has focused on generating practice or assessment questions, in biology exams [32], reading comprehension [3], math practice [31], and programming exercises [25]. As a specific form of question generation, Socratic question generation has gained attention, owing to its effectiveness in improving student learning outcomes by eliciting critical thinking and self-discovery during problem-solving [20].
@@
Socratic questions generation is a complex task that involves mapping out the step-by-step thought process of students during problem-solving, locating the cause of their error, and providing effective questions without revealing the solution. Manually generating Socratic questions can be a cognitively demanding and time-consuming task for instructors. Several recent works proposed to automatically generate Socratic questions using LLMs: In math education, [27] shows that generating a sequence of Socratic sub-questions and prompting students to answer helps them solve math word problems more successfully. In computer science education, [1, 2] releases a dataset on Socratic questions for student code debugging and provides baselines based on LLM prompting and finetuning. In particular, the authors prompt GPT-3.5-turbo and GPT-4 [4] in a chain-of-thought manner [33] to generate Socratic questions. A human study shows that the generated questions can sometimes be invalid in several different ways, including being irrelevant to the problem, repetitive of earlier dialogue turns, or too direct and revealing the solution prematurely, which may hamper students’ learning processes. Since GPT models are proprietary and expensive, the authors also attempt to fine-tune the open-source Flan-T5 model [5]; however, doing so proves to be ineffective due to its insufficient scale and the pertaining procedure used.
@@
In this paper, we propose a method to improve the validity of automatically generating Socratic questions using open-source LLMs. Our method is inspired by recent developments in reinforcement learning with AI feedback (RLAIF) [16]; our method consists of two phases, data augmentation and preference optimization. Specifically, our contributions are as follows:
@@
  • We use the preference information in the dataset, i.e., pairs of valid and invalid Socratic questions, to optimize Llama 2, an open-source LLM, using direct preference optimization (DPO) [22].
  • We show that using the 7B Llama 2 model (with 7B parameters), our best method outperforms existing state-of-the-art methods that rely on larger, proprietary models such as GPT-3.5 and GPT-4, by 7.89 and 3.65 points, respectively, in terms of BERTScore and Rouge-L recall. We also use a series of case studies to illustrate the quality of Socratic questions we can generate and that DPO consistently outperforms supervised fine-tuning (SFT).
2 Background and Related Work
@@ 2.1 Question Generation in Education
In education, question-generation systems are used to create learning materials and problem sets for quizzes and exams. [31] introduces a framework for generating math word problems that incorporates a module for checking the consistency of the word problem generated in terms of the underlying equations that it solves. Our idea of consistency checking of the synthetically generated samples in data augmentation is inspired by theirs. [3] proposes a data augmentation and an over-generate and rank method to fine-tune a language model Flan-T5 [5] to generate questions for reading comprehension. Their data augmentation method prompts a larger LLM to augment the dataset with positive/ valid questions corresponding to a passage in the reading comprehension and then uses this augmented dataset for standard fine-tuning of a smaller open-source LLM. Unlike their work, our data augmentation method involves prompting a larger LLM to generate negative/ invalid questions to create a preference dataset that we use for performing preference optimization on a smaller open-source LLM. In computer science education, recent works show the effectiveness of LLMs like OpenAI Codex and GPT-4 [25, 15] on generating programming exercise questions, code explanations, and test cases. Previous work however does not touch upon guiding students to solve a coding exercise along with maximizing the students’ learning outcomes. [1, 2] introduce a Socratic code debugging dataset, to help a student debug their code along with maximizing the students’ learning outcomes. Their experiments with prompting models like GPT-3.5-turbo, and GPT-4 show that these models tend to hallucinate and produce invalid questions. To address this issue, our work builds upon theirs to fine-tune language models to align the generated questions towards ground-truth human preferences and discourage the models from generating invalid questions.
@@ 2.2 Reward/ Preference Optimization
Fine-tuning language models to align with human preferences has proven to be beneficial in various tasks like machine translation [14], summarization [28, 37], story-telling [37] and instruction fine-tuning [18, 23]. Traditional methods first learn a reward model using a dataset of human preferences and optimize the language model for the downstream task using the rewards obtained from the reward model with reinforcement learning (RL) algorithms such as PPO [26]. There are two drawbacks to this method. First, it is hard to obtain a dataset of human preferences as it is an expensive and sometimes cognitively demanding task. To address this issue procuring rewards from an AI system most commonly an LLM has become a scalable and cheaper alternative [16]. Second, although preference optimization of LLMs using RL algorithms like PPO is effective, it is significantly more challenging and time-consuming than traditional supervised learning as it involves tuning multiple LLMs and sampling rewards in real-time. To address this issue, a new algorithm DPO [22] has been introduced that optimizes a language model to a preference dataset in an RL-free manner by formulating the problem as a binary classification task. DPO significantly reduces the train time and complexity while maintaining similar or even higher performance than traditional PPO methods. In the domain of education, [27] proposes a reward-based method to generate Socratic sub-questions to solve math word problems. Similar to our method they define reward characteristics like fluency, granularity, and answerability to prefer sub-questions that have these desired characteristics. They use REINFORCE [34] a popular RL algorithm to optimize their model by sampling rewards from external systems in real time. Our method is different from theirs as we first prompt an LLM to generate negative/ invalid Socratic questions to construct a preference dataset. We then use this fixed dataset to tune an open-source LLM in an RL-free method, i.e., using DPO which makes the training more stable and less complex. [11] proposes a DPO-based method for fine-tuning LLama 2 [30] for question-answering on a dataset of Piazza posts for an introductory programming course. Their experiments show that DPO consistently outperforms SFT on automatic and human evaluation. However, their process of constructing the preference dataset is straightforward and does not use external LLMs for data augmentation like ours. They create a proxy preference dataset by using the edit history of Piazza posts by preferring the final versions of answers as opposed to the earlier versions.
@@ 3 Problem Definition and Dataset
In this work, we use the dataset for code debugging introduced in  [1, 2]. The dataset is based on didactic conversations between a Student and an Instructor, where the Student is a novice programmer tasked with writing a program for a given problem. The dataset consists of the Student’s buggy code submissions along with a dialogue between the Instructor and the Student, where the Instructor asks Socratic questions in the form of a conversation to help the Student debug their code. The conversation consists of dialogue turns with each Instructor utterance being a collection of several possible “ground-truth” Socratic questions at that dialogue turn. The dataset also contains metadata including the problem statement, the test cases, the bug description, and code fixes to resolve the bug. In total, there are 38 problems with more than 50 different bugs in student solutions, and conversations centered around these buggy codes containing more than 1900 dialogue turns. The dataset is split into two subsets, a train set and a test set which contain 135 and 16 dialogues, respectively, spread across different problems.
@@ 4.1 Data Augmentation
Inspired by methods in RLAIF [16], we augment the dataset with invalid Socratic questions constructed by prompting GPT-4 [4], which provides realistic negative samples for LLM-based question generation methods to train on. We follow the method described in [3] to prompt an LLM to generate synthetic data and employ another instance of the LLM for checking the quality/consistency of the generated synthetic data. Following the definition mentioned in [1], invalid Socratic questions fall into the four following categories:
@@
  • Irrelevant, i.e., questions that are not useful for the student, as they shift focus from the actual bug, which may confuse the student.
@@
  • Repeated, i.e., questions that have already been asked in previous dialogue turns, which are meaningless to the student.
@@
  • Direct, i.e., questions that directly reveal the bug to the student, which do not prompt students to think and may hinder their learning process.
@@
  • Premature, i.e., questions that prompt the student to make code fixes before identifying the bug, which may confuse the student.
@@
To generate invalid questions via an LLM, we construct a few-shot prompt that consists of 1) the definition of the categories as mentioned above and 2) an in-context example for each of the invalid question categories detailed above. Our prompt encourages the model to reason using a chain-of-thought method, by first generating the “reasoning process/logic” behind an invalid question, followed by the question [33]. We generate invalid questions corresponding to all four categories at every dialogue turn where the ground truth is provided.
@@
Following [3, 31], we use a consistency checking step where we prompt GPT-4 to check the consistency of the generated questions to filter out inconsistent questions from the augmented dataset. Inconsistent questions are those that do not belong to any of the invalid categories listed above. We pose the consistency checking step as a classification task where GPT-4 predicts a label for each generated question over six categories, including the four invalid categories and two additional categories: “good” and “incorrect”. Good questions are acceptable Socratic questions at that particular dialogue turn and cannot be used as negative samples. Incorrect questions are unrelated to the problem and the dialogue itself and are often erroneous due to LLM hallucination, which provides little value as easy-to-tell negative samples. We discard all samples that are predicted as “good” or “incorrect”, to get the final set of synthetically generated invalid questions.
@@
Finally, we construct a preference dataset consisting of tuples of valid and invalid Socratic questions. In the preference pairs, valid questions are taken from the ground truth questions in the original dataset, while the invalid questions are generated synthetically as described above. Each valid question from the original dataset is paired with every synthetically generated invalid question of all categories to form the augmented dataset. Since our invalid questions are generated using an LLM potential linguistic or cultural bias related to the pre-training of the LLM might be reflected. However, we hypothesize that this bias would be minimal as Socratic questions are goal-driven, concise, and framed in the second-person perspective directed toward the student.
@@ 4.2 Preference Optimization
In this step, we fine-tune an open-source LLM, Llama 2 [30] for Socratic question generation using DPO [22]. The first step is to perform SFT, i.e., we use the original dataset, $D$, as is to fine-tune LLama 2 for Socratic question generation. For a given conversation in the train set, we first split the dialogue into constituent dialogue turns. The input to LLama 2 is a prompt ($p$) that consists of a systems message that instructs the LLM to generate a Socratic question, the problem metadata, and the current dialogue history (between the Student and the Instructor). The output is the valid Socratic question ($q_{\text{v}}$) corresponding to that dialogue turn in the dataset. In the cases where multiple Socratic questions were given for a dialogue turn, we treat each one as a different output associated with the same input for fine-tuning LLama 2. As shown in Equation 1, the simple SFT step learns a reference policy $\pi_{\text{ref}}$ by minimizing the loss $\mathcal{L}_{SFT}$, which serves as the starting point for preference optimization.
@@
$\mathcal{L}_{SFT}(\pi_{\text{ref}})=$ $-\Huge{\mathbb{E}_{(q_{\text{v}},p)\sim D}}[\log{\pi_{\text{ref}}(q_{\text{v}}|p)}],$ (1)
$\mathcal{L}_{DPO}(\pi_{\theta};\pi_{\text{ref}})=$ $-\Huge{\mathbb{E}_{(q_{\text{v}},q_{\text{iv}},p)\sim D_{P}}}[\log{\sigma(\beta\log{\frac{\pi_{\theta}(q_{\text{v}}|p)}{\pi_{\text{ref}}(q_{\text{v}}|p)}}-\beta\log{\frac{\pi_{\theta}(q_{\text{iv}}|p)}{\pi_{\text{ref}}(q_{\text{iv}}|p)}})}]$ (2)

### v3

@@ Abstract
The Socratic method is a way of guiding students toward solving a problem independently without directly revealing the solution to the problem by asking incremental questions. Although this method has been shown to significantly improve student learning outcomes, it remains a complex labor-intensive task for instructors. Large language models (LLMs) can be used to augment human effort by automatically generating Socratic questions for students. However, existing methods that involve prompting these LLMs sometimes produce invalid outputs, e.g., those that directly reveal the solution to the problem or provide irrelevant or premature questions. To alleviate this problem, inspired by reinforcement learning with AI feedback (RLAIF), we first propose a data augmentation method to enrich existing Socratic questioning datasets with questions that are invalid in specific ways. Also, we propose a method to optimize open-source LLMs such as LLama 2 to prefer ground-truth questions over generated invalid ones, using direct preference optimization (DPO). Our experiments on a Socratic questions dataset for student code debugging show that a DPO-optimized LLama 2-7B model can effectively avoid generating invalid questions, and as a result, outperforms existing state-of-the-art prompting methods [^1] The code for our paper can be found at: https://github.com/umass-ml4ed/socratic-quest-gen.
Improving Socratic Question Generation using Data Augmentation and Preference Optimization
Nischal Ashok Kumar  and Andrew Lan University of Massachusetts Amherst {nashokkumar, andrewlan}@cs.umass.edu
@@ 1 Introduction
Learning based on a conversation that consists of questions and answers, where the student responds to questions posed by a more knowledgeable instructor, has been proven to be effective in teaching students about a particular concept Wood et al. (1976). In particular, Socratic questioning, which refers to a way for the instructor to guide a student to solve a problem (within their zone of proximal development) by asking them questions that promote thinking while not directly revealing the solution Quintana et al. (2018), is a very effective pedagogical method in conversation-based learning and tutoring.
@@
Recent advances in large language models (LLMs) Bubeck et al. (2023) have led to the rapid development of chatbots that promote student learning by automatically generating the instructor’s utterances Dan et al. (2023); Kazemitabaar et al. (2024); Tanwar et al. (2024). One key area of interest in the development of such chatbots is question generation, which can help students solve logical problems in the mathematics and programming domains Al-Hossami et al. (2023); Shridhar et al. (2022). Typically, question generation in educational applications has focused on generating practice or assessment questions, in biology exams Wang et al. (2018), reading comprehension Ashok Kumar et al. (2023), math practice Wang et al. (2021), and programming exercises Sarsa et al. (2022). As a specific form of question generation, Socratic question generation has gained attention, owing to its effectiveness in improving student learning outcomes by eliciting critical thinking and self-discovery during problem-solving Paul and Elder (2007).
@@
Socratic questions generation is a complex task because it involves mapping out the step-by-step thought process of students during problem-solving, locating the cause of their error, and providing effective questions without revealing the solution. Manually generating Socratic questions can be a cognitively demanding and time-consuming task for instructors. Several recent works proposed to automatically generate Socratic questions using LLMs: In math education, Shridhar et al. (2022) shows that generating a sequence of Socratic sub-questions and prompting students to answer helps them solve math word problems more successfully. In computer science education, Al-Hossami et al. (2024, 2023) releases a dataset on Socratic questions for student code debugging and provides baselines based on LLM prompting and finetuning. In particular, the authors prompt GPT-3.5-turbo and GPT-4 Bubeck et al. (2023) in a chain-of-thought manner Wei et al. (2022) to generate Socratic questions. A human study by the authors shows that the generated questions can sometimes be invalid in several different ways, including being irrelevant to the problem, repetitive of earlier dialogue turns, or too direct and revealing the solution prematurely, which may hamper students’ learning processes. Since GPT models are proprietary and expensive, the authors also attempt to fine-tune the open-source Flan-T5 model Chung et al. (2022); however, doing so proves to be ineffective due to its insufficient scale and the pretraining procedure used.
@@
In this paper, we propose a method to improve the validity of automatically generated Socratic questions using open-source LLMs. Our method is inspired by recent developments in reinforcement learning with AI feedback (RLAIF) Lee et al. (2023); our method consists of two phases, data augmentation and preference optimization. Specifically, our contributions are as follows:
@@
  • We use the preference information in the dataset, i.e., pairs of valid and invalid Socratic questions, to optimize Llama 2 Touvron et al. (2023), an open-source LLM, using direct preference optimization (DPO). Rafailov et al. (2023).
  • We show that our method using the Llama 2-7B model outperforms existing state-of-the-art methods that rely on larger, proprietary models such as GPT-3.5 and GPT-4 on the Rouge-L metric and are comparable in terms of BERTScore. We also use a series of case studies to illustrate the quality of Socratic questions we generate and that DPO consistently outperforms supervised fine-tuning (SFT).
2 Related Work
@@ 2.1 Question Generation in Education
In education, question-generation systems are used to create learning materials and problem sets for quizzes and exams. Wang et al. (2021) introduces a framework for generating math word problems that incorporates a module for checking the consistency of the word problem generated in terms of the underlying equations that it solves. Our idea of checking the consistency of the synthetically generated samples in data augmentation is inspired by theirs. Ashok Kumar et al. (2023) proposes a data augmentation and an over-generate and rank method to fine-tune a language model Flan-T5 Chung et al. (2022) to generate questions for reading comprehension. Their data augmentation method prompts a larger LLM to augment the dataset with valid questions (positive examples) corresponding to a passage in the reading comprehension and then uses this augmented dataset for standard fine-tuning of a smaller open-source LLM. Unlike their work, our data augmentation method involves prompting a larger LLM to generate invalid questions (negative examples) to create a preference dataset that we use for performing preference optimization on a smaller open-source LLM. In computer science education, recent works show the effectiveness of LLMs like OpenAI Codex and GPT-4 Sarsa et al. (2022); Kumar and Lan (2024) on generating programming exercise questions, code explanations, and test cases. Al-Hossami et al. (2024, 2023) introduce a Socratic code debugging dataset, to help a student debug their code along with maximizing the students’ learning outcomes. Their experiments with prompting models like GPT-3.5-turbo, and GPT-4 show that these models tend to hallucinate and produce invalid questions. To address this issue, our work builds upon theirs to fine-tune language models to align the generated questions towards ground-truth human preferences and discourage the models from generating invalid questions.
@@ 2.2 Reward/ Preference Optimization
Fine-tuning language models to align with human preferences has proven to be beneficial in various natural language processing tasks Kreutzer et al. (2018); Stiennon et al. (2020); Ziegler et al. (2019); Ouyang et al. (2022). Traditional methods first learn a reward model using a dataset of human preferences and optimize the language model for the downstream task using the rewards obtained from the reward model with reinforcement learning (RL) algorithms such as PPO Schulman et al. (2017). There are two drawbacks to this method. First, it is hard to obtain a dataset of human preferences as it is an expensive and sometimes cognitively demanding task. To address this issue, RLAIF procures rewards from an AI system, such as an LLM, and has become a scalable and cheaper alternative Lee et al. (2023). Second, although preference optimization of LLMs using RL algorithms like PPO is effective, it is significantly more challenging and time-consuming than traditional supervised learning as it involves tuning multiple LLMs and sampling rewards in real time. To address this issue, the DPO method Rafailov et al. (2023) optimizes a language model to a preference dataset in an RL-free manner by formulating the problem as a binary classification task.
In the domain of education, Shridhar et al. (2022) proposes a reward-based method to generate Socratic sub-questions to solve math word problems. Similar to our method they define reward characteristics like fluency, granularity, and answerability to prefer sub-questions that have these desired characteristics. They use REINFORCE Williams (1992) a popular RL algorithm to optimize their model by sampling rewards from external systems in real time. Our method is different from theirs as we first prompt an LLM to generate invalid Socratic questions (negative examples) to construct a preference dataset. We then use this fixed dataset to tune an open-source LLM in an RL-free method, i.e., using DPO which makes the training more stable and less complex. Hicke et al. (2023) proposes a DPO-based method for fine-tuning LLama 2 Touvron et al. (2023) for question-answering on a dataset of Piazza posts for an introductory programming course. They create a proxy preference dataset by using the edit history of Piazza posts by preferring the final versions of answers as opposed to the earlier versions. However, the setting of their work is different from ours as we focus on Socratic question generation and propose a method to create the preference dataset using data augmentation. Scarlatos et al. (2024) propose a method to perform DPO on LLama 2 for the task of feedback generation to help students solve mathematics word problems. To create preference pairs they prompt LLMs like Codex Chen et al. (2021) and GPT-3.5 turbo to generate bad feedback and rate the feedback based on a pre-defined rubric using GPT-4. Our problem setting is different from theirs as we focus on the programming education domain and for our task the LLM needs to provide a series of step-by-step feedback in the form of a dialogue-based interaction through Socratic questions instead of just providing the feedback once for a given problem.
@@ 3 Problem Definition and Dataset
In this work, we use the dataset for code debugging introduced in  Al-Hossami et al. (2024, 2023). The dataset is based on didactic conversations between a Student and an Instructor, where the Student is a novice programmer tasked with writing a program for a given problem. The dataset consists of the Student’s buggy code submissions along with a dialogue between the Instructor and the Student, where the Instructor asks Socratic questions in the form of a conversation to help the Student debug their code. The conversation consists of dialogue turns with each Instructor utterance being a collection of several possible “ground-truth” Socratic questions at that dialogue turn. The dataset also contains metadata including the problem statement, the test cases, the bug description, and code fixes to resolve the bug. In total, there are 38 problems with more than 50 different bugs in student solutions, and conversations centered around these buggy codes containing more than 1900 dialogue turns. The dataset is split into two subsets, a train set and a test set which contain 135 and 16 dialogues, respectively, spread across different problems.
@@ 4.1 Data Augmentation
Inspired by methods in RLAIF Lee et al. (2023), we augment the dataset with invalid Socratic questions constructed by prompting GPT-4 Bubeck et al. (2023), which provides realistic negative samples for LLM-based question generation methods to train on. We follow the method described in Ashok Kumar et al. (2023) to prompt an LLM to generate synthetic data and employ another instance of the LLM for checking the quality/consistency of the generated synthetic data. Following the definition mentioned in Al-Hossami et al. (2024), invalid Socratic questions fall into the four following categories:
@@
  • Irrelevant questions that are not useful for the student, as they shift focus from the actual bug, which may confuse the student.
@@
  • Repeated questions that have already been asked in previous dialogue turns, which are meaningless to the student.
@@
  • Direct questions that directly reveal the bug to the student, which do not prompt students to think and may hinder their learning process.
@@
  • Premature questions which prompt the student to make code fixes before identifying the bug, which may confuse the student.
@@
To generate invalid questions via an LLM, we construct a few-shot prompt that consists of 1) the definition of the categories as mentioned above and 2) an in-context example for each of the invalid question categories detailed above. Our prompt encourages the model to reason using a chain-of-thought method, by first generating the “reasoning process/logic” behind an invalid question, followed by the question Wei et al. (2022). We generate invalid questions corresponding to all four categories at every dialogue turn where the ground truth is provided.
@@
Following Ashok Kumar et al. (2023); Wang et al. (2021), we use a consistency checking step where we prompt GPT-4 to check the consistency of the generated questions to filter out inconsistent questions from the augmented dataset. Inconsistent questions are those that do not belong to any of the invalid categories listed above. We pose the consistency checking step as a classification task where GPT-4 predicts a label for each generated question over six categories, including the four invalid categories and two additional categories: “good” and “incorrect”. Good questions are acceptable Socratic questions at that particular dialogue turn and cannot be used as negative samples. Incorrect questions are unrelated to the problem and the dialogue itself and are often erroneous due to LLM hallucination, which provides little value as easy-to-tell negative samples. To maintain high data quality of our preference dataset, we discard all samples that are predicted as “good” or “incorrect”, to get the final set of synthetically generated invalid questions.
@@
Finally, we construct a preference dataset consisting of 2500 tuples of valid and invalid Socratic questions. In the preference pairs, valid questions are taken from the ground truth questions in the original dataset, while the invalid questions are generated synthetically as described above. Each valid question from the original dataset is paired with every synthetically generated invalid question of all categories to form the augmented dataset.
@@ 4.2 Preference Optimization
In this step, we fine-tune an open-source LLM, Llama 2 Touvron et al. (2023) for Socratic question generation using DPO Rafailov et al. (2023). The first step is to perform SFT, i.e., we use the original dataset, $D$, as is to fine-tune LLama 2 for Socratic question generation. For a given conversation in the train set, we first split the dialogue into constituent dialogue turns. The input to LLama 2 is a prompt ($p$) that consists of a systems message that instructs the LLM to generate a Socratic question, the problem metadata, and the current dialogue history (between the Student and the Instructor). The output is the valid Socratic question ($q_{\text{v}}$) corresponding to that dialogue turn in the dataset. In the cases where multiple Socratic questions were given for a dialogue turn, we treat each one as a different output associated with the same input for fine-tuning LLama 2. As shown in Equation 1, the simple SFT step learns a reference policy $\pi_{\text{ref}}$ by minimizing the loss $\mathcal{L}_{SFT}$, which serves as the starting point for preference optimization.
@@
$\mathcal{L}_{\text{SFT}}(\pi_{\text{ref}})=-\mathbb{E}_{(q_{\text{v}},p)\sim D}[\log{\pi_{\text{ref}}(q_{\text{v}}|p)}]$ (1)
$\mathcal{L}_{DPO}(\pi_{\theta};\pi_{\text{ref}})=\\ -\mathbb{E}_{(q_{\text{v}},q_{\text{iv}},p)\sim D_{P}}\bigg{[}\log{\sigma(\beta\log{\frac{\pi_{\theta}(q_{\text{v}}|p)}{\pi_{\text{ref}}(q_{\text{v}}|p)}}}\\ -\beta\log{\frac{\pi_{\theta}(q_{\text{iv}}|p)}{\pi_{\text{ref}}(q_{\text{iv}}|p)}})\bigg{]}$ (2)